OpenStudy (anonymous):

Evaluate 4h-2, when h = 9

OpenStudy (anonymous):

h = 9 4(9)-2 = (4x9)-2 (4x9)-2 = ?

OpenStudy (anonymous):

put h=9 in 4h-2 u will get the answer as 36-2=34